# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ Abstract
Index Terms:
Beyond-diagonal reconfigurable intelligent surfaces, modes, reciprocal and non-reciprocal architecture designs, reconfigurable impedance unitary property.
@@ I Introduction
The sixth generation (6G) networks are driven by three important characteristics linked to human lifestyle and social changes in the near future: High-fidelity holographic society; connectivity for all things; and time sensitive/time engineered applications [1, 2, 3]. Consequently, 6G networks should be human-centric with the aim of achieving high security and privacy, seamless connectivity between devices, and real-time communications with extremely low latency [3]. To achieve these stringent requirements, several promising 6G candidate solutions have been proposed, such as advanced multiple access techniques that could better utilize resources [4], extremely large-scale multiple input multiple output (XL-MIMO) that could significantly improve the spectral efficiency [5], multifunctional platforms that could simultaneously support wireless communications, sensing, and power transfer [6, 7], and reconfigurable intelligent surfaces (RISs) that provide wave manipulation flexibility to reconfigure the wireless propagation environment and enable new processing capabilities in the electromagnetic domain [8, 9, 10, 11].
@@ I-A Background and Motivation
Among various promising 6G candidate solutions, RIS has gained significant attention in recent years due to its capability to change the way of treating wireless propagation environments: from adapting to them by sophisticated signal processing strategies to manipulating them with low cost, negligible thermal noise, and low power consumption [8]. RIS is a generic term that is also referred to as intelligent reflecting surface (IRS) [9], intelligent surface (IS) [12], and programmable/dynamic metasurface [8]. RIS can be physically fabricated using different implementations, such as using antenna arrays with tunable components and using metasurfaces, while the former way gains more attention in the wireless society since it has a more straightforward and physics-consistent model. Broadly speaking, an RIS is a planar surface consisting of a large number of tunable components. These tunable components can be realized by varactors and/or positive intrinsic negative (PIN) diodes, each of which is able to induce a controllable change of phase shift and/or amplitude to the incident signal. Benefiting from its shape and reconfigurable property, RIS can be flexibly deployed in complex wireless propagation environments, e.g., being attached on high buildings, to bypass obstacles and increase/decrease the directivity of waves according to specific requirements. RISs were primarily proposed as nearly passive devices, in the sense that they are not capable of amplifying incident waves and only minimal power is used to control the surface [9]. This property brings various practical advantages for RIS implementation, such as being free of power-hungry RF chains, power amplifiers, and thermal noise.
@@
Inspired by the above advantages, the research on RIS has grown exponentially in the past few years, covering a wide range of areas that include but are not limited to beamforming design [13, 14], channel estimation [15, 16], hardware impairments analysis [17, 18, 19], and implementation and prototyping [11, 20]. To fully compensate for the huge path loss induced by RIS, the concept of active RIS, named due to the introduction of reflection-type power amplifiers, has been further proposed to provide significantly enhanced performance gain at the expense of affordable power consumption and cost [21]. Meanwhile, the industry has also achieved progress on demonstrating and testing RISs. The test mainly originated from NTT DOCOMO, a Japanese network operator, which demonstrated a fifth generation (5G) mobile system using a metasurface reflectarray operating in 28 GHz in November 2018 and conducted a trail of a transparent dynamic metasurface for 5G radio signals in January 2020, followed by a re-conduction in 2021 [8]. More recently, RIS has been used in a 5G commercial frequency range 2 (FR2) band network that operates beyond 24 GHz, showing that the indoor coverage can be significantly expanded and the throughput improvement can reach 227.69% [22]. In support of active academia and industrial progress on RIS, key organizations have started to consider the integration of RIS into future commercial networks. For example, RIS-related standardization has been kicked off in the China Communications Standards Association and the FuTURE Mobile Communication Forum from China [23]. For another example, within the European Telecommunications Standards Institute (ETSI), an industry specific group on RISs has been proposed and approved in June 2021, and launched in September 2021 [23]. In the first phase from 2021 to 2023, the use cases, deployment scenarios, technical challenges, models, and evaluation methods of applying RIS have been discussed [24]. The second phase has also been kicked off in September 2023 with the focus on more specific practical considerations for RIS, more detailed use cases, deployment scenarios, transmission and reception schemes, and more functional RIS supporting transmission, reflection, sensing, and computing [12].
@@
$\mathbf{\Theta}=\left[\begin{matrix}?&&&\\ &?&&\\ &&\ddots&\\ &&&?\end{matrix}\right]\Rightarrow\mathbf{\Theta}=\left[\begin{matrix}?&?&\ldots&?\\ ?&?&\ldots&?\\ \vdots&&\ddots&\vdots\\ ?&?&\cdots&?\end{matrix}\right].$ (1)
The answer is yes, with the emerging technology, namely beyond-diagonal (BD) RIS [25].
@@ I-B BD-RIS
BD-RIS refers to a novel and general framework for RIS whose scattering matrix $\mathbf{\Theta}$ is not limited to be diagonal, hence not only diagonal entries, but also off-diagonal entries which have been forced to zeros in D-RIS can be flexibly tuned to manipulate waves [25, 26]. This is realized by inter-connecting (part of) elements by additional reconfigurable components as illustrated in Fig. 1(b), hence allowing waves absorbed by one element to flow through other elements and generating nonzero and tunable off-diagonal entries in $\mathbf{\Theta}$. This brings extra degrees of freedom (DoF) for RIS from the following two perspectives.
@@
  • Element Arrangement Flexibility: The way of arranging elements in D-RIS is limited due to the lack of inter-element interactions. Specifically, since elements in D-RIS are typically designed to cover half of space, they are often arranged as a 2D planar array, leading to practical location constraints for transmitters and receivers, as illustrated in Figs. 2(a) and 2(b). By contrast, BD-RIS enables more flexible element arrangements thanks to inter-element connections that allow signals to flow from one direction to another, as illustrated in Figs. 2(c) and 2(d). BD-RIS with diverse element arrangements thus provides opportunities to achieve larger wireless coverage and denser connectivity.
@@
Due to the above extra DoF provided by flexible inter-element reactions in BD-RIS, it is expected that BD-RIS will be able to achieve smarter wireless transmissions with higher quality, denser connectivity, and wider coverage. Fig. 3 shows an envisioned future application scenario supported by BD-RIS. For instance, BD-RIS can be particularly useful to extend the coverage in joint indoor and outdoor communications in Fig. 3. Properly deploying BD-RIS with flexible element arrangements can also help serving cell-edge users especially when there are obstacles between the base station and users. Another interesting application is to enable flexible and scalable integrated access and backhaul (IAB) [27]. BD-RIS can be flexibly incorporated into real and complex environments to assist and enhance wireless backhauling between macrocells and picocells and wireless access between picocells and users. In addition to wireless communications, with proper power levels, suitable deployments, and locations, BD-RIS can be potentially used for higher-efficiency power relay and wireless power transfer (WPT) [7]. Besides, BD-RIS can also be leveraged to boost wireless sensing, especially in complicated propagation conditions, such as vehicle networks, where strong line-of-sight (LoS) links are not available between the radar and targets. Not limited to stand-alone functions (wireless communications, power transfer, and sensing), BD-RIS can support more advanced dual-functional systems, such as simultaneous wireless information and power transfer (SWIPT) [28] and integrated sensing and communications (ISAC) [6], or multi-functional integrated communications, sensing, and power transfer systems [29]. Not limited to being used to realize smart radio environment, BD-RIS can also be integrated with transmitters/receivers and work as an auxiliary antenna array to enable RIS-aided or stacked surface-aided massive MIMO. Beyond the above, BD-RIS, as a comprehensive upgrade of D-RIS, is readily applicable to all D-RIS enabled scenarios while providing more flexible deployment, smarter wave manipulation and better performance, and wider coverage.
@@
[25] Overview
The first magazine to provide an overview
of BD-RIS and to discuss its benefits,
challenges, and potential applications
[30, 31]
Briefly revisit the fundamentals of BD-RIS
and non-terrestrial networks (NTNs), and
outline the application of BD-RIS in NTNs
[32]
Highlight the advantages of BD-RIS for
Internet of Things (IoT) applications and
present technical challenges
[33] Survey
Provide a survey of BD-RIS techniques and
state-of-the art research, and point out
future research directions
@@ I-C Contributions and Organization
The appealing benefits and potential of BD-RIS have spurred active research. Specifically, a few overviews [25, 30, 31, 32] and survey [33] papers have appeared, focusing more on potential application scenarios and future research directions of BD-RIS, as summarized in Table I. However, there is a lack of a comprehensive tutorial that pedagogically explains the basics and fundamental tools to understand BD-RIS. The main goal of this paper is thus to provide the first tutorial on BD-RIS including the fundamentals from a rigorous microwave engineering perspective, the state-of-the-art signal processing techniques, the hardware impairments, and the benefits, applications, technical challenges, and future research directions of BD-RIS. This further leads to the following contributions.
@@
Organization: The rest of this paper is organized as follows. Section II provides a toy example of BD-RIS-aided SISO systems to briefly explain the rationale and benefits of BD-RIS. Section III explains the basic modeling, mode analysis, and various architecture designs for BD-RIS. Section IV summarizes the key techniques for BD-RIS optimization and channel estimation. Section V introduces hardware impairments in BD-RIS and illustrates their impacts on system performance. Section VI discusses the benefits of BD-RIS supported by analytical and numerical results, and summarizes emerging applications. Section VII points out challenges which should be tackled to facilitate the use of BD-RIS in future 6G networks, and sheds light on possible future research directions. Finally, Section VIII concludes this paper. The outline of this paper is illustrated in Table II.
@@ C. Contributions and Organization
A. Modeling and Classification
B. Mode Analysis
C. Architecture Design: Reciprocal Architectures
D. Architecture Design: Non-Reciprocal Architectures
E. Unified Modes and Architectures
@@ C. Channel Estimation
Section V. BD-RIS with Hardware Impairments
@@ D. Mutual Coupling Effect
Section VI. Benefits and Applications of BD-RIS
A. Benefits
B. Emerging Applications
Section VII. Challenges and Future Research Directions of BD-RIS
@@ E. New BD-RIS Architectures
Section VIII. Conclusion
Appendix. Preliminaries on Multi-Port Network Analysis
A. Impedance, Admittance, and Scattering Parameters
B. Reciprocal and Lossless Networks
Notations: Boldface lower- and upper-case letters indicate column vectors and matrices, respectively. $(\cdot)^{\mathsf{T}}$, $(\cdot)^{*}$, $(\cdot)^{\mathsf{H}}$, and $(\cdot)^{-1}$ denote the transpose, conjugate, conjugate-transpose, and inverse operations, respectively. $\mathbb{C}$, $\mathbb{R}$, and $\mathbb{Z}$ denote the sets of complex, real, and integer numbers, respectively. $\mathbb{E}\{\cdot\}$ denotes the statistical expectation. ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$ denote the real and imaginary parts of complex numbers, respectively. ($\angle$($\cdot$)∠ ( ⋅ ) denotes the angle of complex numbers. $\mathsf{blkdiag}(\cdot)$ represents a block-diagonal matrix and $\mathsf{diag}(\cdot)$ represents a diagonal matrix. ||$\cdot$|| ⋅ |, $\|\cdot\|_{2}$, and $\|\cdot\|_{\mathsf{F}}$ denote the absolute value of a scalar, the $\ell_{2}$ norm of a vector, and the Frobenius norm of a matrix, respectively. $\jmath=\sqrt{-1}$ denotes the imaginary unit. $\otimes$ denotes the Kronecker product. $\mathsf{vec}(\cdot)$ denotes the vectorization operation. $\mathsf{tr}(\cdot)$ denotes the trace of a matrix. $\det(\cdot)$ denotes the determinant of a square matrix. $\mathbf{I}_{M}$ denotes an $M\times M$ identity matrix. $\mathbf{0}_{M\times N}$ denotes an $M\times N$ all-zero matrix. $\mathbf{A}\preceq\mathbf{B}$ (or $\mathbf{A}\succeq\mathbf{B}$) indicates that $\mathbf{B}-\mathbf{A}$ (or $\mathbf{A}-\mathbf{B}$) is positive semi-definite. $a\sim\mathcal{CN}(0,\sigma^{2})$ characterizes the circular symmetric complex Gaussian distribution. $[\mathbf{a}]_{i}$ denotes the $i$-the entry of $\mathbf{a}$. $[\mathbf{A}]_{i:i^{\prime},j:j^{\prime}}$ extracts the $i$-th to $i^{\prime}$-th rows and the $j$-th to $j^{\prime}$-th columns of $\mathbf{A}$.
@@ II BD-RIS-Aided SISO: A Toy Example
where $e^{\jmath\theta_{m}}$ is the reflection coefficient of the $m$-th RIS element. Note that we have implicitly assumed the D-RIS to be lossless in (3) by considering reflection coefficients with unit magnitude. Thus, the channel gain maximization problem in a D-RIS-aided SISO system writes as
$\underset{\boldsymbol{\Theta}}{\mathsf{\mathrm{max}}}$ $\left|\mathbf{h}_{RI}\boldsymbol{\Theta}\mathbf{h}_{IT}\right|^{2}$ (4)
$\mathsf{\mathrm{s.t.}}$ $\boldsymbol{\Theta}=\mathrm{diag}\left(e^{\jmath\theta_{1}},e^{\jmath\theta_{2}},\ldots,e^{\jmath\theta_{M}}\right).$
It is well-known that a global optimal solution to (4) is obtained by co-phasing all $M$ paths reflected by the $M$ RIS elements, such that they interfere constructively at the receiver. Mathematically, such a solution is given in closed form as
$\boldsymbol{\Theta}^{\star}$ $=\mathrm{diag}\left(e^{\jmath\theta_{1}^{\star}},e^{\jmath\theta_{2}^{\star}},\ldots,e^{\jmath\theta_{M}^{\star}}\right),$ (5)
$\theta_{m}^{\star}$ $=-\angle\left(\left[\mathbf{h}_{RI}\right]_{m}\left[\mathbf{h}_{IT}\right]_{m}\right),\>\forall m\in\mathcal{M},$
where =$\mathcal{M}=\{1,\ldots,M\}$. This allows us to obtain a channel gain
$\rho^{\mathsf{D-RIS}}=\left(\sum_{m=1}^{M}\left|\left[\mathbf{h}_{RI}\right]_{m}\left[\mathbf{h}_{IT}\right]_{m}\right|\right)^{2},$ (6)
as it can be verified by substituting (5) in (2).
@@
The matrix $\boldsymbol{\Theta}$ is commonly referred to as the “phase shift matrix” of the RIS, because conventionally it is a diagonal matrix containing phase shifts in its diagonal, as given by (3). Nevertheless, it has been shown through microwave theory that, more rigorously speaking, $\boldsymbol{\Theta}$ is the so-called “scattering matrix” of the RIS. In microwave theory, the scattering matrix is a matrix that can be used to characterize a linear microwave network linking the reflected waves to the incident waves at its ports. Intuitively speaking, the scattering matrix can be seen as the multidimensional generalization of the concept of reflection coefficient. As a lossless 1-port microwave network is characterized by a reflection coefficient having unit magnitude, a lossless multi-port microwave network is characterized by a unitary scattering matrix [34].
@@
Since a lossless microwave network has a unitary scattering matrix, a lossless RIS is in principle allowed to have any unitary matrix $\boldsymbol{\Theta}$, i.e., $\boldsymbol{\Theta}^{\mathsf{H}}\boldsymbol{\Theta}=\mathbf{I}_{M}$ or, equivalently, $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathsf{H}}=\mathbf{I}_{M}$, and is not limited to the diagonal constraint in (3). We refer to these RIS architectures, which are not limited to having a diagonal scattering matrix $\boldsymbol{\Theta}$, as BD-RIS. Remarkably, given that the unitary constraint $\boldsymbol{\Theta}^{\mathsf{H}}\boldsymbol{\Theta}=\mathbf{I}_{M}$ is a generalization of the constraint in (3), BD-RIS is a generalization of D-RIS and includes D-RIS as a special case. The channel gain maximization problem in the presence of a BD-RIS can be written as
$\underset{\boldsymbol{\Theta}}{\mathsf{\mathrm{max}}}$ $\left|\mathbf{h}_{RI}\boldsymbol{\Theta}\mathbf{h}_{IT}\right|^{2}$ (7)
$\mathsf{\mathrm{s.t.}}$ $\boldsymbol{\Theta}^{\mathsf{H}}\boldsymbol{\Theta}=\mathbf{I}_{M},$
which can be solved through a global optimal solution as follows. To globally solve (7), we first present an upper bound on the objective function, and then we present a solution for $\boldsymbol{\Theta}$ proved to achieve that upper bound. First, following the sub-multiplicity property of the spectral norm, we obtain that the channel gain in (7) is upper bounded by
@@
based on the Cauchy–Schwarz inequality and the constraint $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$. Second, it is possible to find a solution for $\boldsymbol{\Theta}$ to achieve that upper bound in closed form. Consider the matrix $\mathbf{V}_{RI}\in\mathbb{C}^{M\times M}$ as a unitary matrix containing the right singular vectors of $\mathbf{h}_{RI}$ in its columns. In other words, the first column of $\mathbf{V}_{RI}$ is $[\mathbf{V}_{RI}]_{:,1}=\frac{\mathbf{h}_{RI}^{\mathsf{H}}}{\|\mathbf{h}_{RI}\|_{2}}$ and the other columns of $\mathbf{V}_{RI}$ are unit-norm vectors mutually orthogonal and orthogonal to $[\mathbf{V}_{RI}]_{:,1}$. Similarly, consider the matrix $\mathbf{U}_{IT}\in\mathbb{C}^{M\times M}$ as a unitary matrix containing the left singular vectors of $\mathbf{h}_{IT}$ in its columns. In other words, the first column of $\mathbf{U}_{IT}$ is $[\mathbf{U}_{IT}]_{:,1}=\frac{\mathbf{h}_{IT}}{\|\mathbf{h}_{IT}\|_{2}}$ and the other columns of $\mathbf{U}_{IT}$ are unit-norm vectors mutually orthogonal and orthogonal to $[\mathbf{U}_{IT}]_{:,1}$. Given $\mathbf{V}_{RI}$ and $\mathbf{U}_{IT}$, it is easy to see that a unitary matrix $\boldsymbol{\Theta}$ globally solving problem (7) is
$\boldsymbol{\Theta}^{\star}=\mathbf{V}_{RI}\mathbf{U}_{IT}^{\mathsf{H}},$ (9)
which allows to achieve the channel gain upper bound in (8).

### v2

@@ I Introduction
The sixth generation (6G) networks are driven by three important characteristics linked to human lifestyle and social changes in the near future: High-fidelity holographic society; connectivity for all things; and time sensitive/time engineered applications [tataria20216g, wang2023road, dang2020should]. Consequently, 6G networks should be human-centric with the aim of achieving high security and privacy, seamless connectivity between devices, and real-time communications with extremely low latency [dang2020should]. To achieve these stringent requirements, several promising 6G candidate solutions have been proposed, such as advanced multiple access techniques that could better utilize resources [clerckx2024multiple], extremely large-scale multiple input multiple output (MIMO) that can significantly improve the spectral efficiency [lu2024tutorial], multifunctional platforms that could simultaneously support wireless communications, sensing, and power transfer [liu2022integrated, clerckx2021wireless], and reconfigurable intelligent surfaces (RISs) that provide wave manipulation flexibility to reconfigure the wireless propagation environment and enable new processing capabilities in the electromagnetic domain [di2020smart, wu2025intelligent, kaina2014shaping, tang2020wireless].
@@ I-A Background and Motivation
Among various promising 6G candidate solutions, RIS has gained significant attention in recent years due to its capability to change the way of treating wireless propagation environments: from adapting to them by sophisticated signal processing strategies to manipulating them with low cost, negligible thermal noise, and low power consumption [di2020smart]. RIS is a generic term that is also referred to as intelligent reflecting surface (IRS) [wu2025intelligent], intelligent surface (IS) [wu2024intelligent], and programmable/dynamic metasurface [di2020smart]. RIS can be physically fabricated using different implementations, such as using antenna arrays with tunable components and using metasurfaces, while the former way gains more attention in the wireless society since it has a more straightforward and physics-consistent model. Broadly speaking, an RIS is a planar surface consisting of numerous tunable elements, each of which is able to induce a controllable change of phase shift and/or amplitude to the incident signal. Benefiting from its shape and reconfigurable property, RIS can be flexibly deployed in complex wireless propagation environments, e.g., being attached on high buildings, to bypass obstacles and increase/decrease the directivity of waves according to specific requirements. RISs were primarily proposed as nearly passive devices, in the sense that they are not capable of amplifying incident waves and only minimal power is used to control the surface [wu2025intelligent]. This property brings various practical advantages for RIS implementation, such as being free of power-hungry RF chains, power amplifiers, and thermal noise.
@@
Inspired by the above advantages, the research on RIS has grown exponentially in the past few years, covering a wide range of areas that include but are not limited to beamforming design [guo2020weighted, wu2019intelligent], channel estimation [swindlehurst2022channel, zheng2022survey], hardware impairments analysis [abeywickrama2020intelligent, gradoni2021end, li2021intelligent], and implementation and prototyping [tang2020wireless, araghi2022reconfigurable]. To fully compensate for the huge path loss induced by RIS, the concept of active RIS, named due to the introduction of reflection-type power amplifiers, has been further proposed to provide significantly enhanced performance gain at the expense of affordable power consumption and cost [zhang2022active]. Meanwhile, the industry has also achieved progress on demonstrating and testing RISs. The test mainly originated from NTT DOCOMO, a Japanese network operator, which demonstrated a fifth generation (5G) mobile system using a metasurface reflectarray operating in 28 GHz in 2018 and conducted a trail of a transparent dynamic metasurface for 5G radio signals in 2020, followed by a re-conduction in 2021 [di2020smart]. More recently, RIS has been used in a 5G commercial frequency range 2 band network that operates beyond 24 GHz, showing significant performance improvement in the indoor coverage and throughput [yang2023beyond]. In support of active academia and industrial progress on RIS, key organizations have started to consider the integration of RIS into future commercial networks. For example, RIS-related standardization has been kicked off in the China Communications Standards Association and the FuTURE Mobile Communication Forum from China [liu2022path]. For another example, within the European Telecommunications Standards Institute, an industry specific group on RISs has been proposed and approved in June 2021, and launched in September 2021 [liu2022path]. The technical details and practical considerations for applying RISs have been thoroughly discussed during 2021-2023 [etsi, wu2024intelligent].
@@
[…]
The answer is yes, with the emerging technology, namely beyond-diagonal (BD) RIS [li2023reconfigurable].
@@ I-B BD-RIS
BD-RIS refers to a novel and general framework for RIS whose scattering matrix $\mathbf{\Theta}$ is not limited to be diagonal, hence not only diagonal entries, but also off-diagonal entries which have been forced to zeros in D-RIS can be flexibly tuned to manipulate waves [li2023reconfigurable, shen2021]. This is realized by inter-connecting (part of) elements by additional reconfigurable components as illustrated in Fig. 1(b), hence allowing waves absorbed by one element to flow through other elements and generating nonzero and tunable off-diagonal entries in $\mathbf{\Theta}$. This brings extra degrees of freedom (DoF) for RIS from the following two perspectives.
@@
  • Element Arrangement Flexibility: In D-RIS, the arrangement of the elements is constrained by the absence of inter-element interconnections. Specifically, since elements in D-RIS are typically designed to cover half of space, they are often arranged as a 2D planar array, leading to practical location constraints for transmitters and receivers, as illustrated in Figs. 2(a) and 2(b). By contrast, BD-RIS enables more flexible element arrangements thanks to inter-element connections that allow signals to flow from one direction to another, as illustrated in Figs. 2(c) and 2(d). BD-RIS with diverse element arrangements thus provides opportunities to achieve larger wireless coverage and denser connectivity.
@@
Due to the above extra DoF provided by flexible inter-element reactions in BD-RIS, it is expected that BD-RIS will be able to achieve smarter wireless transmissions with higher quality, denser connectivity, and wider coverage. Fig. 3 shows an envisioned future application scenario supported by BD-RIS. For instance, BD-RIS can be particularly useful to extend the coverage in joint indoor and outdoor communications in Fig. 3. Properly deploying BD-RIS with flexible element arrangements can also help serving cell-edge users especially when there are obstacles between the base station and users. Another interesting application is to enable flexible and scalable integrated access and backhaul [madapatha2020integrated]. BD-RIS can be flexibly incorporated into real and complex environments to assist and enhance wireless backhauling between macrocells and picocells and wireless access between picocells and users. In addition to wireless communications, with proper power levels, suitable deployments, and locations, BD-RIS can be potentially used for higher-efficiency power relay and wireless power transfer (WPT) [clerckx2021wireless]. Besides, BD-RIS can also be leveraged to boost wireless sensing, especially in complicated propagation conditions, such as vehicle networks, where strong line-of-sight (LoS) links are not available between the radar and targets. Not limited to stand-alone functions (wireless communications, power transfer, and sensing), BD-RIS can support more advanced dual-functional systems, such as simultaneous wireless information and power transfer (SWIPT) [perera2017simultaneous] and integrated sensing and communications (ISAC) [liu2022integrated, hua2023secure], or multi-functional integrated communications, sensing, and power transfer systems [zhou2024integrating]. Not limited to being used to realize smart radio environment, BD-RIS can also be integrated with transmitters/receivers and work as an auxiliary antenna array to enable RIS-aided or stacked surface-aided massive MIMO. Beyond the above, BD-RIS, as a comprehensive upgrade of D-RIS, is readily applicable to all D-RIS enabled scenarios while providing more flexible deployment, smarter wave manipulation and better performance, and wider coverage.
@@
[li2023reconfigurable] Overview
The first magazine to provide a high-level
overview of the concept of BD-RIS
without technical details
[khan2024beyond, khan2024integration, khan2025beyond]
Briefly revisit the fundamentals of BD-RIS,
non-terrestrial networks (NTNs), and Internet
of Things (IoT), and outline the application
of BD-RIS in NTNs and IoT Networks
@@ I-C Contributions and Organization
The appealing benefits and potential of BD-RIS have spurred active research. Specifically, a few overviews [li2023reconfigurable, khan2024beyond, khan2024integration, khan2025beyond] have appeared, focusing either on high-level illustrations to BD-RIS or specific applications of BD-RIS, as summarized in Table I. However, there is a lack of a comprehensive tutorial that pedagogically explains the basics and fundamental tools to understand BD-RIS. The main goal of this paper is thus to provide the first tutorial on BD-RIS including the fundamentals from a rigorous microwave engineering perspective, the state-of-the-art signal processing techniques, the hardware impairments, and the benefits, applications, technical challenges, and future research directions of BD-RIS. This further leads to the following contributions.
@@
Organization: The rest of this paper is organized as follows. Section II provides a toy example of BD-RIS-aided SISO systems to briefly explain the rationale and benefits of BD-RIS. Section III explains the basic modeling, mode analysis, and various architecture designs for BD-RIS. Section IV summarizes the key techniques for BD-RIS optimization and channel estimation. Section VI introduces hardware impairments in BD-RIS and illustrates their impacts on system performance. Section V discusses the benefits of BD-RIS supported by analytical and numerical results, and summarizes emerging applications. Section VIII points out challenges which should be tackled to facilitate the use of BD-RIS in future 6G networks, and sheds light on possible future research directions. Finally, Section IX concludes this paper. The outline of this paper is illustrated in Table II.
@@ C. Contributions and Organization
A. Preliminaries on Multi-Port Network Analysis
B. Modeling and Classification
C. Mode Analysis
D. Architecture Design: Reciprocal Architectures
E. Architecture Design: Non-Reciprocal Architectures
F. Unified Modes and Architectures
@@ C. Channel Estimation
Section V. Benefits of BD-RIS
A. Boosting Received Power and Rates
B. Enabling Low-Complexity Architectures with High Performance
C. Enabling Flexible Modes with Highly-Directional Wireless Coverage
D. Providing Orders of Magnitude Gains in Distributed Deployments
E. Enabling Simultaneously Optimal Transmissions for Uplink and
      Downlink with Non-Reciprocal Architectures
F. Providing Enhanced Gains in Dual-Polarized Systems
Section VI. BD-RIS with Hardware Impairments
@@ D. Mutual Coupling Effect
Section VII. Applications of BD-RIS
A. BD-RIS for Communications
B. BD-RIS for Sensing and ISAC
C. BD-RIS for WPT and SWIPT
D. BD-RIS with Other Techniques and Systems
Section VIII. Challenges and Future Research Directions of BD-RIS
@@ E. New BD-RIS Architectures
Section IX. Conclusion
Notations: Boldface lower- and upper-case letters indicate column vectors and matrices, respectively. $(\cdot)^{\mathsf{T}}$, $(\cdot)^{*}$, $(\cdot)^{\mathsf{H}}$, and $(\cdot)^{-1}$ denote the transpose, conjugate, conjugate-transpose, and inverse operations, respectively. $\mathbb{C}$, $\mathbb{R}$, and $\mathbb{Z}$ denote the sets of complex, real, and integer numbers, respectively. The superscript of $(\cdot)^{M\times N}$ identifies the size of a matrix. $\mathbb{E}\{\cdot\}$ denotes the statistical expectation. $\Re\{\cdot\}$ and $\Im\{\cdot\}$ denote the real and imaginary parts of complex numbers, respectively. $\angle(\cdot)$ denotes the angle of complex numbers. $\mathsf{blkdiag}(\cdot)$ represents a block-diagonal matrix and $\mathsf{diag}(\cdot)$ represents a diagonal matrix. $|\cdot|$, $\|\cdot\|_{2}$, and $\|\cdot\|_{\mathsf{F}}$ denote the absolute value of a scalar, the $\ell_{2}$ norm of a vector, and the Frobenius norm of a matrix, respectively. $\jmath=\sqrt{-1}$ denotes the imaginary unit. $\otimes$ denotes the Kronecker product. $\mathsf{vec}(\cdot)$ denotes the vectorization operation. $\mathsf{tr}(\cdot)$ denotes the trace of a matrix. $\det(\cdot)$ denotes the determinant of a square matrix. $\mathbf{I}_{M}$ denotes an $M\times M$ identity matrix. $\mathbf{0}_{M\times N}$ denotes an $M\times N$ all-zero matrix. $\mathbf{A}\preceq\mathbf{B}$ (or $\mathbf{A}\succeq\mathbf{B}$) indicates that $\mathbf{B}-\mathbf{A}$ (or $\mathbf{A}-\mathbf{B}$) is positive semi-definite. $a\sim\mathcal{CN}(0,\sigma^{2})$ characterizes the circular symmetric complex Gaussian distribution. $\mathbf{a}]_{i}$ denotes the $i$-the entry of $\mathbf{a}$. $\mathbf{A}]_{i:i^{\prime},j:j^{\prime}}$ extracts the $i$-th to $i^{\prime}$-th rows and the $j$-th to $j^{\prime}$-th columns of $\mathbf{A}$.
@@ II BD-RIS-Aided SISO: A Toy Example
where $e^{\jmath\theta_{m}}$ is the reflection coefficient of the $m$-th RIS element. Note that we have implicitly assumed the D-RIS to be lossless in (3) by considering reflection coefficients with unit magnitude. The detailed derivation for obtaining such a diagonal phase shift matrix at D-RIS will soon be shown in Sections III-B and III-D. We can now write the channel gain as $|h|^{2}=|\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{h}_{IT}|^{2}=|\sum_{m=1}^{M}[\mathbf{h}_{RI}]_{m}e^{\jmath\theta_{m}}[\mathbf{h}_{IT}]_{m}|^{2}$. This allows us to obtain an optimal phase shift of the $m$-th element of RIS as $\theta_{m}^{\star}=-\angle([\mathbf{h}_{RI}]_{m}[\mathbf{h}_{IT}]_{m})$, yielding a maximum channel gain
$\rho^{\mathsf{D-RIS}}=\left(\sum_{m=1}^{M}\left|\left[\mathbf{h}_{RI}\right]_{m}\left[\mathbf{h}_{IT}\right]_{m}\right|\right)^{2}.$ (4)
@@
The matrix $\boldsymbol{\Theta}$ is commonly referred to as the “phase shift matrix” of the RIS, because conventionally it is a diagonal matrix containing phase shifts in its diagonal, as given by (3). Nevertheless, it has been shown through microwave theory that, more rigorously speaking, $\boldsymbol{\Theta}$ is the so-called “scattering matrix” of the RIS. In microwave theory, the scattering matrix is a matrix that can be used to characterize a linear microwave network linking the reflected waves to the incident waves at its ports. Intuitively speaking, the scattering matrix can be seen as the multidimensional generalization of the concept of reflection coefficient. As a lossless 1-port microwave network is characterized by a reflection coefficient having unit magnitude, a lossless multi-port microwave network is characterized by a unitary scattering matrix [pozar2021microwave].
@@
Since a lossless microwave network has a unitary scattering matrix, a lossless RIS is in principle allowed to have any unitary matrix $\boldsymbol{\Theta}$, i.e., $\boldsymbol{\Theta}^{\mathsf{H}}\boldsymbol{\Theta}=\mathbf{I}_{M}$ or, equivalently, $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathsf{H}}=\mathbf{I}_{M}$, and is not limited to the diagonal constraint in (3). $\mathbf{\Theta}$ in this case can have various expressions based on specific circuit designs, each of which will be explained in detail in Sections III-D and III-E. We refer to these RIS architectures, which are not limited to having a diagonal scattering matrix $\boldsymbol{\Theta}$, as BD-RIS. Remarkably, given that the unitary constraint $\boldsymbol{\Theta}^{\mathsf{H}}\boldsymbol{\Theta}=\mathbf{I}_{M}$ is a generalization of the constraint in (3), BD-RIS is a generalization of D-RIS and includes D-RIS as a special case. In this case, following the sub-multiplicity property of the spectral norm, we obtain that the channel gain $|h|^{2}$ is upper bounded by
@@
based on the Cauchy-Schwarz inequality and the constraint $\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}$.